Adriana Corporation manufactures football equipment. In planning for next year, the managers want to understand the relation between activity and overhead costs. Discussions with the plant supervisor suggest that overhead seems to vary with labor-hours, machine-hours, or both. The following data were collected from last year's operations:

Month Labor-Hours Machine-Hours Overhead Costs
1 730 1,354 $ 102,672
2 720 1,400 103,841
3 670 1,525 109,998
4 735 1,459 108,270
5 770 1,584 116,165
6 745 1,588 114,486
7 730 1,400 106,963
8 720 1,309 102,087
9 715 1,459 106,404
10 795 1,551 113,024
11 680 1,298 104,420
12 700 1,608 117,595


Required:

a.

a. Use the high-low method to estimate the fixed and variable portions of overhead costs based on machine-hours. (Round "Variable cost" answer to 2 decimal places.)


b. Managers expect the plant to operate at a monthly average of 1,600 machine-hours next year. What are the estimated monthly overhead costs, assuming no inflation? (Round "Variable cost" answer to 2 decimal places.)
Answer #1

a.

Variable cost per machine hour = (Highest activity cost – Lowest activity cost)/(Highest activity machine hour – Lowest activity machine hour)

= (117,595 - 104,420)/(1,608 - 1,298)

= 13,175/310

= $42.5

Fixed cost = Highest activity cost – (Variable cost per machine hour x Highest activity machine hour)

= 117,595 - (42.5 x 1,608)

= 117,595 - 68,340

= $49,255

b.

Total cost when 1,600 machine hours used = Fixed cost + (Variable cost per machine hour x 1,600 )

= 49,255 + (42.5 x 1,600)

= 49,255 + 68,000

= $117,255

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.